# Understanding Long Videos in One Multimodal Language Model Pass

### v1

@@ 5.2 Robotics Domain Action Recognition
In this section, we evaluate the generalization capabilities of the proposed methodologies across various subsets of the Open X-Embodiment robotics dataset [40]. The Open X-Embodiment dataset is an extensive collection of visuomotor robotics datasets, encompassing a wide range of tasks and environments. It is designed to facilitate research in visuomotor control, providing rich sensory inputs and motor outputs for training and testing robotic systems. However, the videos are usually taken in a controlled environment and they do not always contain meaningful objects, which makes the samples in the dataset usually out of general video distribution (See Fig. 4).
For our analysis, we specifically select subsets of the dataset that contain expert episodes accompanied by corresponding language instructions and adapt them into video classification datasets. We treat each trajectory within the dataset as a video clip, with its associated language instruction serving as the video caption (classification label). For each episode, the model is tasked with identifying the correct language instruction from a set of five options, considering a video clip uniformly downsampled to 8 frames. The incorrect options are randomly chosen from the dataset to ensure a diverse and challenging selection. In instances where the datasets have multiple cameras for each episode, we treat the videos captured by each camera as distinct datasets.
Figure 4: Examples frames of different videos from EgoSchema dataset (top) and examples from Open X-Embodiment dataset (bottom). Compared to top row, images below appear out of distribution given the controlled environment and robot movement.
@@
Our results presented in Tab. 3 indicate clear performance improvements for our proposed MVU over both random selection and our competitive single frame baseline. The purpose of this experiment is to evaluate the generality of our approach to video domains different from everyday natural videos. We take these promising results in Tab. 3 to indicate the strong generality of our framework.
5.3 Modality Constrained Evaluation
We evaluate the two modality-constrained variants of our approach, Just LLM and Single Frame VLM (details in Sec. 4.4) and summarize these findings in Tab. 4. We uncover surprisingly strong performance of both variants on two long-video understanding benchmarks. Note how these approaches use no video-specific information to generate predictions.
We highlight how our best Just LLM variant achieves performance significantly higher than random selection (+25.8% on EgoSchema-S / +20.1% on NextQA-T) using zero visual information. This indicates the large portion of questions in existing video-QA benchmarks that can be answered correctly purely using world knowledge. We also highlight our single frame variant performing on par with some existing state-of-the-art (gray). In particular, for EgoSchema-S we outperform [83] which uses information extracted from 180 frames per video incurring an inference cost over 100 times higher than ours. In light of these findings, we argue that long video understanding approaches in particular must focus on learning information beyond what a single frame baseline can achieve.
We also take these results as an indication of the clear inference speed-up from Likelihood Selection (more specific ablations in Sec. 5.4).
Table 4: Modality Constrained Long Video Evaluation: We evaluate our proposed Vid-LLM and Vid-VLM setup, on two video question-answering datasets that specifically focus on longer videos. Number of samples in each dataset inside brackets. EgoSchema-S is public subset of EgoSchema [35] and NextQA-T is the test set of NextQA dataset [65]. LS backbone is the pre-trained model used for likelihood selection. We report Top-1 Accuracy (standard evaluation protocol) and inference time per sample (single video-question pair) when running on two NVIDIA RTX A5000 24GB GPUs. Recent state-of-the-art work [83] using multi-frame inputs (gray) is also reported for comparison. indicates the use of additional LLMs/VLMs for intermediate processing.
Method LS Backbone EgoSchema-S (500) NextQA-T (8564)
Acc (%) Time (s) Acc (%) Time (s)
Random - 20.0 - 20.0 -
Just LLM (ours) Llama-2-7b-Chat 17.4 0.89 28.8 1.17
Just LLM (ours) Gemma-7b-IT 45.8 2.38 39.3 2.60
Just LLM (ours) Mistral-7B-Instruct 45.8 0.41 40.1 0.55
SF-VLM (our) LLaVA-v1.5-13B 55.8 1.89 51.2 2.03
LLoVi [83] Mistral-7B-Instruct 50.8 381 54.3 207
@@ 5.4 Ablations
We systematically dissect our overall framework to establish the usefulness of each individual component of our method. We first ablate our overall MVU framework and report these results in Tab. 5. Our evaluations indicate clear performance boosts from both our proposed frame selection strategy and each of our object-centric modality information.
We next perform ablations on our likelihood selection strategy and report these results in Tab. 6. Our results indicate clear performance boosts due to likelihood selection (LS) in terms of both accuracy and inference speed. In the alternate case (no LS), a standard post-generation selection [83] (i.e. generate an answer and match against ground-truth selection) is utilized. While our inference speed boosts against this setup can be directly attributed to our single pass aspect, we hypothesize that accuracy gains arise as a result of reduced LLM output hallucination and deviations from expected output formats, that are commonly observed in other approaches [18].
We also conduct an ablation on our object motion trajectory modality inputs focussed on the fine-grained action recognition dataset, Something-Something-v2 [15] (SSv2). As elaborated in Sec. 5, SSv2 is a dataset focused on motion-based category discrimination, providing an ideal evaluation to measure the usefulness of our object motion trajectory modality. We benchmark on a subset of this dataset following [31] and report these results in Tab. 7. Our results while exceeding their performance also indicate the clear performance gains obtained when injecting the object motion trajectory modality into our MVU framework.
Table 5: Ablation on object-centric modality information. We report accuracy (top-1 %) on the 500 sample subset of EgoSchema. In the table header, VI and FS stand for visual inputs and frame selection respectively. GOI, OSL, OMT refer to our object-centric modalities (see Sec. 4.2). Each component of our approach, FS and the integration of separate modality information, results in clear performance improvements. Our overall framework, MVU (our-full), achieves the best performance.
Method VI FS GOI OSL OMT EgoSchema-S
Just LLM (ours) 45.8
Single Frame VLM (ours) 55.8
@@
Table 6: Ablation on Likelihood Selection. Results indicate performance improvements in terms of both top-1 accuracy (%) and inference time (s) when utilizing our likelihood selection approach.
Method LS Accuracy Time
MVU (ours) 56.4 8.65
MVU (ours) 60.3 2.42
Table 7: Ablation on OMT modality. Results in the motion-based SSv2 dataset indicate the usefulness of our OMT modality.
Method OMT Accuracy
Random - 0.6
CLIP [44] - 4.0
MAXI [31] - 6.4
MVU (ours) 3.6
MVU (ours) 7.2
@@ 6 Conclusion
In this work, we present a multimodal video understanding framework, termed MVU, that achieves state-of-the-art performance on complex video understanding tasks. In particular, the evaluation of long-video question answering and fine-grained action recognition (language-tied zero-shot) demonstrates the strong performance of our MVU framework. We propose an efficient technique for LLM-based answer choice selection and extend this for frame selection highlighting its generality. We separate video-specific information into three object-centric modalities, propose a language-based fusion of this multimodal information, and explore their individual contributions to our end performance.
Revisiting our original motivation on, “what we can learn from videos, beyond scene context understood from a single natural image”, in this work we also presented two modality-constrained variants of our approach that uncover surprising insights relevant to this question. We achieve strong results on long-video understanding benchmarks using no video-specific data, and build over that baseline to showcase the additional performance gains achievable through injecting video-specific information.
@@
Our method utilizes multiple pretrained models, all of which are open-source with model weights freely available. We use the versions of these models hosted on HuggingFace https://huggingface.co/ for all our experiments. We discuss all steps in our proposed algorithms in detail while also releasing relevant code. All evaluations we perform are on public datasets accessible by all (some behind evaluation servers to prevent test set contamination). Our code is hosted at https://github.com/kahnchana/mvu.
Ethics Statement
We present a framework specific for long-video understanding fusing individual video modalities through language and also explore efficient inference strategies for LLM-based methods. Although we believe our algorithms and framework introduce no additional negative biases (especially given how we do not train any new models), we acknowledge our use of existing models that are known to contain certain biases (from the datasets they are trained on). However, in our work, we consider some of these biases (i.e. world knowledge) as actual strengths that should be given careful consideration when evaluating such LLM-based frameworks even on complex downstream tasks. Additionally, we highlight the interpretability of our language-based information fusion, which allows explicit observation of the types of biases present, allowing easier alterations for cases that require the absence of certain biases.
Acknowledgements
This work was supported by the SUNY Empire Innovation Program (EIP). We thank the Robotics Lab at SBU, especially Jongwoo Park, for helpful discussions.
Contributions
KR led the project coming up with the initial idea, building the preliminary codebase, and performing most experimentation. XL was responsible for extending the framework to the robotics domain, contributed ideas for extracting OSL and OMT modality information, and constructed all figures in the paper. KK ran preliminary experiments on verifying our likelihood selection strategy, performed multiple long-video question-answering experiments, and discussed all aspects of the project. MR organized and supervised the entire project, set the research direction, and discussed all aspects of the project.
@@ Acknowledgements
Appendix 0.A Prompting and Template Operations
In Sec. 4.2 and Sec. 4.3, we utilize 3 distinct prompts and fusion templates for generating joint textual inputs to be processed by the LLM. The 3 distinct prompt categories correspond to our Global Object Information ($x_{\mathrm{GOI}}$), Object Spatial Location ($x_{\mathrm{OSL}}$), and Object Motion Trajectory ($x_{\mathrm{OMT}}$) modalities. We first describe our exact templates as Python pseudo-code in Table A.1.
@@ Appendix 0.A Prompting and Template Operations
Appendix 0.B Frame Selection Queries
As described in Sec. 4.1, we utilize our likelihood selection strategy for selecting frames optimal to solve a selected task. In this setting, we fix the textual input questions $x_{t}$ and possible target $y$ (only one target) and alter the input visual tokens based on different frames $x_{v}^{l}$ ($l$ is the index of video frame), in order to obtain per-frame likelihood measures. The prompt template at presented at Table A.3
User <image to select $x_{v}^{l}$> Consider task: task_question. How useful is this frame to answer the given question?   Assistant This frame is highly useful and contains relevant information
Table A.3: Prompt template for frame selection.
We hypothesize that frames with more relevant information will maximize the likelihood of our target text and utilize our promising results (see performance gains from frame selection in Table 5) as validation.
Appendix 0.C Tokenization in LLMs
Most modern LLMs utilize Byte-Pair Encoding (BPE) tokenization [56] to convert natural language into discrete tokens that are mapped to vocabulary embeddings. This process is learned from language itself and the resulting tokenization may sometimes break complete words into pieces (e.g. example $\rightarrow$ ex-am-ple). Given our utilization of logits extracted from an LLM forward pass, we note that each logit corresponds to a single token, which may at times be the embedding of some meaningless word piece. However, our calculation of a joint likelihood across a sequence of tokens ensures a meaningful process, which is validated by our strong results.
Appendix 0.D Details on Baselines
@@
In Sec. 5.1, we evaluate the performance on long-video understanding tasks using work in [83, 61] are two baselines for comparison. However, both these methods utilize closed-source, proprietary LLMs (i.e. GPT-4) with parameter counts on the order of trillions (over 100X our model size) deeming their unfair comparisons. In the case of [83], we replicate their method (using their open-source repository and pre-trained models following [22]) utilizing an open-source LLM of comparable parameter count as our framework. For [61], the authors directly report results for a variant with a similar parameter count as ours. We utilize these evaluations as our point of comparison.

### v2

@@ 5.2 Robotics Domain Action Recognition
In this section, we investigate generalization capabilities of our proposed MVU by evaluating across various datasets from robotics domain Open X-Embodiment (Open-X-Embodiment-Collaboration et al., 2023), following a QnA style formulation of the dataset. See Appendix D for more details on this dataset. We highlight the visual differences of this data in Figure 4.
Figure 4: Data Visualization: Example video frames from EgoSchema (top) vs OpenX (bottom) datasets. Robotics domain videos (bottom) appear out of distribution given their controlled environment and robot movements.
@@
Our results on these robotics datasets are presented in Table 4 and indicate clear performance improvements for our proposed MVU over the baseline from Zhang et al. (2023a). The purpose of this experiment is to evaluate the generality of our approach to video domains different from everyday natural videos. We take these promising results to indicate the strong generality of our framework.
Furthermore, we note how our modality constrained variants do not perform significantly better than random on these robotics domain tasks (details in Appendix E). We attribute this to the significant domain shift in terms of the world of operation in this domain (i.e. robotics tasks tend to involve controlled environments very different to what humans face on an everyday basis).
Table 4: OpenX Detailed Results: We report accuracy (%) for the VideoQA formulation of Open X-Embodiment benchmark (Open-X-Embodiment-Collaboration et al., 2023). Our proposed MVU achieves clear improvements over random selection and LLoVi baseline (Zhang et al., 2023a). In the table header, Obs. (observation), size, CC (class count) stand for the camera used, number of videos, and number of unique language instructions per dataset, respectively. In the observation column, T stands for third-person view (stationary camera that does not move with robot), while F denotes first-person view where camera is mounted on moving robot. Note that total is average weighted by dataset size.
Dataset Obs. Size CC Random Baseline MVU
ASU TableTop Manipulation T 110 83 13.6 19.1 20.9
Berkeley MVP Data F 480 6 20 26.0 33.1
Berkeley RPT Data F 908 4 24.6 23.1 26.2
CMU Play Fusion T 576 44 20.3 34.0 35.6
CMU Stretch T 135 5 23 18.5 24.4
Furniture Bench T 5100 9 20.2 24.8 26.4
Furniture Bench F 5100 9 20.2 22.6 24.9
CMU Franka Pick-Insert Data T 631 7 18.7 19.3 21.2
CMU Franka Pick-Insert Data F 631 7 23.1 57.8 49.3
Imperial F Cam T 170 17 20 22.9 24.1
Imperial F Cam F 170 17 23.5 20.6 24.7
USC Jaco Play T 1085 89 21.8 26.4 30.6
USC Jaco Play F 1085 89 19.4 28.6 32.4
NYU ROT T 14 12 21.4 57.1 57.1
Roboturk T 1959 3 34.7 43.0 44.2
Stanford HYDRA T 570 3 35.1 54.7 68.2
Stanford HYDRA F 570 3 31.2 45.3 48.9
Freiburg Franka Play F 3603 406 20.4 32.2 31.6
Freiburg Franka Play T 3603 406 19.7 21.8 24.0
LSMO Dataset T 50 2 34.0 68.0 72.0
UCSD Kitchen T 150 8 19.3 32.0 32.7
Austin VIOLA T 150 3 26.7 32.7 33.3
Austin VIOLA F 150 3 30.0 33.3 34.0
Total - 27000 - 22.1 28.5 30.4
@@ 5.3 Ablations
In this section, we systematically dissect our overall MVU framework to establish the usefulness of each of its individual component. We first ablate our three different information modalities and report these results in Table 5. Our evaluations indicate clear performance boosts from each of our object-centric information modalities.
Table 5: MVU Ablation: We report accuracy (%) on public subset of EgoSchema (ES-S). In table header, VI stand for visual inputs and GOI, OSL, OMT refer to our object-centric information modalities (see Section 4.2). Each information modality results in clear performance improvements with our full MVU achieving best performance.
Method VI GOI OSL OMT ES-S
Just-LLM (ours) 45.8
SF-VLM (ours) 55.8
@@
We next perform ablations on our adaptation of likelihood selection strategy for video QnA tasks using Ego-Schema subset (ES-S) and Next-QA test-set (NQA-T) . These results reported in Table 7 indicate clear performance boosts due to our adaptation of likelihood selection (LS). When removing LS, standard generation (i.e. generate an answer and match against ground-truth selection following Zhang et al. (2023a)) is utilized with our MVU framework. We also report naive adaptation of LS following Robinson et al. (2023) where the choice options are directly used, highlighting the importance of our prompting techniques. We also note the accuracy gains obtained through LS, and attribute these to reduced LLM output hallucination and deviations from expected output formats, that are commonly observed with iterative generation (Hanu et al., 2023).
We next ablate our overall framework against the existing work Zhang et al. (2023a) by replacing our MVU object-centric information pipeline with the frame description approach in Zhang et al. (2023a). We construct a setup identical to our framework except for the inputs to our final stage VLM replaced with frame level descriptions. These results reported in Table 7 indicate the clear significance and improvement of our proposed object-centric information pipeline over simple frame descriptions. The 8 frame variant is the same speed comparison as MVU uses captioner only on 8 frames. Our MVU outperforms both that and the slower 16 frame baseline. We also note the performance drop in the baseline when increasing the number of frames from 16 to 180. While consistent with observations in prior works for long-video tasks (Mangalam et al., 2023), we attribute this drop to decreased signal to noise ratio with introducing of additional frame descriptions. This further highlights the importance of selecting useful information from video frames and we reiterate how the object-centric information in our MVU framework serves this purpose.
Table 6: Likelihood Selection (LS) Ablation: Results indicate clear improvements in both accuracy (%) and inference time (s) with our adaptation of likelihood selection for video tasks.
Method LS ES-S NQA-T Time
Generation 56.4 55.3 12.7
LS-Naive 58.2 35.8 2.42
LS-MVU (ours) 60.3 55.4 2.42
Table 7: Baseline Ablation: We replace information input to final stage VLM with frame descriptions following Zhang et al. (2023a). Accuracy (%) on public subset of EgoSchema (ES-S). Time in seconds (s).
Method Frames ES-S Time
Baseline 180 55.4 207
Baseline 8 55.8 2.38
Baseline 16 56.2 4.72
MVU (ours) 16 60.3 2.42
@@ 6 Conclusion
In this work, we present a multimodal video understanding framework, termed MVU, that achieves state-of-the-art performance on complex video understanding tasks. In particular, evaluations on long-video question answering and robotics domain question answering demonstrate the strong performance of our MVU framework as well as its generality. We also adapt likelihood selection for efficient LLM-based answer choice selection, separate video-specific information into three object-centric modalities, demonstrate automated extraction of such information using off-the-shelf vision tools, and propose language-based fusion of this multimodal information.
We also presented two modality-constrained baselines that uncover surprising insights relevant to LLM based video QnA which serves as a basis for our subsequent MVU framework. Additionally, these results highlight the need for careful evaluation of LLM-based video QnA approaches. Revisiting our original motivation on, “what we can learn from videos, beyond scene context understood from a single natural image”, in this work our two modality-constrained variants uncover surprising insights relevant to this question. We first achieve strong results on long-video understanding benchmarks using no video-specific data, and build over that baseline to showcase the additional performance gains achievable through injecting video-specific information.
@@
Our method utilizes multiple pretrained models, all of which are open-source with model weights freely available. We use the versions of these models hosted on HuggingFace https://huggingface.co for all our experiments. We discuss all steps in our proposed algorithms in detail while also releasing relevant code. All evaluations we perform are on public datasets accessible by all (some behind evaluation servers to prevent test set contamination). Our code will also be released publicly.
@@
Appendix A Prompting and Template Operations
In Section 4.2 and Section 4.3, we utilize 3 distinct prompts and fusion templates for generating joint textual inputs to be processed by the LLM. The 3 distinct prompt categories correspond to our Global Object Information ($x_{\mathrm{GOI}}$), Object Spatial Location ($x_{\mathrm{OSL}}$), and Object Motion Trajectory ($x_{\mathrm{OMT}}$) modalities. We first describe our exact templates as Python pseudo-code in Table A.1.
@@ Appendix A Prompting and Template Operations
Appendix B Details on Pretrained Models and Datasets
We describe in detail the pretrained models used to construct our framework as well as the multiple datasets used in evaluating our framework.
Models: Our framework utilizes three distinct off-the-shelf models for its various operations, namely a) an LLM / VLM for likelihood selection (including visual frame selection and textual answer selection), b) a generative VLM for extracting object list from a frame, and c) an open-vocabulary detector for object localization. We use LLaVA-v1.5-13B (Liu et al., 2023a) for likelihood selection, frame selection, and frame object list generation. For object localization, we use OWL-ViT-B/32 (Minderer et al., 2022). Unless explicitly specified, we use the above setup in all our experiments. Variants of our framework uses LLMs Llama-2-7b-Chat, Gemma-7b-IT, and Mistral-7B-Instruct (default) for likelihood selection. Apart from these off-the-shelf models, our framework involves zero additional training. We also reiterate that no components of our framework undergo any form of video-level training.
Datasets: We use multiple video datasets for evaluation under question-answering or n-way classification settings. For video question answering, we select two datasets focused on long-form videos: EgoSchema (Mangalam et al., 2023), NExT-QA (Xiao et al., 2021). EgoSchema is a long-form ego-centric video question-answering benchmark, consisting of a 500-video public subset (EgoSchema-S) and a full 5000+ video evaluation set (EgoSchema-F) accessed only through evaluation servers. This dataset spans over 250 hours and is specially constructed to ensure that questions require awareness of a longer temporal window for correctly answering (Mangalam et al., 2023). Example images of EgoSchema are shown in Figure 4. NExT-QA similarly contains long-form videos with a focus on requiring causal & temporal action reasoning as well as common scene comprehension for correctly answering. It contains a validation set (NExT-QA-V) of 4996 video-questions pairs and a test set (NExT-QA-T) of 8564 video-question pairs. We also use a series of robotics datasets from the Open X-Embodiment robotics dataset (Open-X-Embodiment-Collaboration et al., 2023) for video question answering in a different domain (more detail in Section 5.2). In only one of our ablations aimed at analysing the motion understanding aspect of our framework, we utilize a fine-grained action recognition dataset, Something-Something-v2 (Goyal et al., 2017), that contains 174 action categories focusing on object motions by replacing object category nouns with ‘something’ in textual descriptions of each action category.
Appendix C Details on Baselines
@@
In Section 5.1, we evaluate the performance on long-video understanding tasks using work in Zhang et al. (2023a) and Wang et al. (2023) as two baselines for comparison. However, both these methods utilize closed-source, proprietary LLMs (i.e. GPT-4) with parameter counts on the order of trillions (over 100X our model size) deeming their direct comparisons unfair. In the case of Zhang et al. (2023a), we replicate their method (using their open-source repository and pre-trained models following Kahatapitiya et al. (2024)) utilizing an open-source LLM of comparable parameter count as our framework. For Wang et al. (2023), the authors directly report results for a variant with a similar parameter count as ours. We utilize these evaluations as our point of comparison.
Appendix D Robotics Domain Dataset Details
The Open X-Embodiment dataset is an extensive collection of visuomotor robotics datasets, encompassing a wide range of tasks and environments. It is designed to facilitate research in visuomotor control, providing rich sensory inputs and motor outputs for training and testing robotic systems. However, the videos are usually taken in a controlled environment and they do not always contain meaningful objects, which makes the samples in the dataset usually out of general video distribution (See Figure 4).
For our analysis, we specifically select datasets within this collection that contain expert episodes accompanied by corresponding language instructions and adapt them into video classification datasets. We treat each trajectory within the dataset as a video clip, with its associated language instruction serving as the video caption (classification label). For each episode, the model is tasked with identifying the correct language instruction from a set of five options, considering a video clip uniformly downsampled to 8 frames. The incorrect options are randomly chosen from the dataset to ensure a diverse and challenging selection. In instances where the datasets have multiple cameras for each episode, we treat the videos captured by each camera as distinct datasets.
Appendix E Discussion on Modality Constrained Evaluation
We evaluate the two modality-constrained variants of our approach, Just-LLM and Single-Frame-VLM (details in Section 3.3) and summarize these findings in Table 1. We uncover surprisingly strong performance of both variants on two long-video understanding benchmarks. Note how these approaches use no video-specific information to generate predictions.
We highlight how our best Just-LLM variant achieves performance significantly higher than random selection (+25.8% on EgoSchema-S / +20.1% on NextQA-T) using zero visual information. This indicates the large portion of questions in existing video-QnA benchmarks that can be answered correctly purely using world knowledge. We also highlight our single frame variant performing on par with some existing state-of-the-art (gray). In particular, for EgoSchema-S we outperform Zhang et al. (2023a) which uses information extracted from 180 frames per video incurring an inference cost over 100 times higher than ours. In light of these findings, we argue that long video understanding approaches in particular must focus on learning information beyond what a single frame baseline can achieve.
We also evaluate these same modality-constrained variants on robotics domains tasks and report these results in Table A.3. In contrast to the results on standard long-video QnA benchmarks, the robotics domains results are more aligned with intuition: the no-visual input Just-LLM performs on par with random and the Single-Frame-VLM marginally outperforms random selection.
We attribute this difference in performance to the nature of robotics domain tasks. They tend to involve controlled environments with often naive, meaningless tasks purely for robot testing purposes. These may not necessarily align with human commonsense or other constraints dependent on knowledge of our world. Therein, the clear ability of LLMs to solve general everyday video tasks (e.g. EgoSchema, NextQA performance in Table 1) using its world knowledge may not be applicable to robotics domain tasks. Utilizing different domain benchmarks, in particular robotics tasks, provides a much more representative evaluation of LLM based video QnA approaches.
Table A.3: Modality Constrained Variants on Robotics Domain: We evaluate our modality constrained baselines on the robotics domain tasks and report accuracy (%). Note that a weighted sum over multiple tasks is reported here (similar to Table 4). Note the minimal increase over random for the variants in contrast to generic video benchmarks.
Method Visual Frames Accuracy
Random - - 22.1
Just-LLM - 21.9
SF-VLM 1 23.5
MVU 16 30.4
Appendix F Likelihood Selection
In this section, we present the prompts and templates used to adapt likelihood selection inference (Robinson et al., 2023) to our video QnA tasks. Our experimentation shows significantly higher sensitivity (to prompt style) of LLM performance on QnA tasks when using like likelihood selection in comparison to sequential text generation (consistent with findings in Robinson et al. (2023)). We evaluate a series of different prompt templates on the EgoSchema and Next-QA dataset to discover optimal combinations. The best prompt templates used in our final framework are presented in Table A.4 as Python pseudo-code. For Next-QA in particular, the average zero-shot accuracy could vary from 35% to 55% with slight variations of the prompt templates.
Our optimal prompt templates for the standard video QnA tasks also generalized to our robotics domain QnA tasks. Nevertheless, we highlight the possibility of needing some prompt template tuning when applying our framework to different domains. We also note that while our prompt selection process was guided by heuristics and experimentation, there may be other similar prompts performing equally well or surpassing our optimal selection.
F.1 Implementation Details
We revisit the generation process of autoregressive LLMs and their visual extensions (VLMs). They commonly use iterative prediction of next tokens conditioned on prior outputs to generate complete natural language outputs. Such a generation process is usually modeled as sampling from a conditional likelihood shown as Equation 2, where $\hat{y}^{j}$ stands for the $j^{\mathrm{th}}$ token in a textual sequence $\hat{y}$ autoregressively generated by the model.
$P(\hat{y}|x_{t})$ $=\prod_{j}P(\hat{y}^{j}|\hat{y}^{1,...,j-1},x_{t})$ (2)
$P(y_{i}|x_{t})$ $=\prod_{j}P(y^{j}_{i}|y^{1,...,j-1}_{i},x_{t})$ (3)
Assume a perfect LLM, intuitively when $y_{i}$ is a proper answer to the question $x_{t}$ (say $y_{i}=y_{g}$), the conditional likelihood $P(y_{i}|x_{t})$ should be larger than any other $P(y_{w}|x_{t})$ where $y_{w}$ is a wrong answer to question $x_{t}$. In fact, modern LLMs are trained with a similar objective (Radford & Narasimhan, 2018). Motivated by this standard LLM training objective, we estimate the relative numerical scales of conditional likelihood on different answers $P(y_{i}|x_{t})$ using a cross-entropy error $e_{i}$, given their equivalence (negative log-likelihood and multiclass cross-entropy, see Section 4.3.4 in Bishop (2006)). We calculate $e_{i}$ with a single forward pass of LLM without detokenization and the selection can be made by simply picking up the answer with the lowest error, equivalent to the highest conditional likelihood among all the answers.
[…]
Table A.4: Likelihood Selection Sample Prompt Templates. Variables qs and prompt_list refer to per sample question and choice list respectively.
This sets the ground for Likelihood Selection, also referred to as Cloze Promting in Robinson et al. (2023), first illustrated with a toy example in Figure 2, where the task is vanilla question-answering with only textual context and the model takes one question $x_{t}$ as well as $M=5$ candidate answers $y_{1,...,5}$. To find the best answer, we simply concatenate the question with each candidate independently ($s_{i}=\mathrm{concat}\left(x_{t},y_{i}\right)$ ) and pad them into a batch $\{s_{1,...,5}\}$. Then the LLM takes the batch of five samples with causal attention masks and performs one inference forward pass, resulting in five shifted logits $\{p_{1,...,5}\}$. Next, we shift the input sequence $s_{i}$ to align the logits $p_{i}$ and calculate the average cross-entropy error only on tokens of $y_{i}$ Finally, the answer with the smallest $e_{i}$ will be picked up as the best answer. The method can be formulated as in Equation 5 using equivalence of negative log-likelihood to cross-entropy in Equation 4. Here $n_{i}$ stands for the token sequence length of $y_{i}$ and $p_{i}^{j}$ stands for logits of the $j^{\mathrm{th}}$ token in $p_{i}=V(\mathrm{concat}\left(x_{t},y_{i}\right))$ with logits limited to only those of $y_{i}$.
In summary, Likelihood Selection performs one single forward pass to extract the network logit outputs, calculates error ($e_{i}$) on each choice, and selects the choice with the lowest error. Note that our method does not utilize any iterative autoregressive generation using the LLM. This results in considerable speed-ups for inference time. We also obtain the additional advantages of avoiding LLM hallucinations and deviations from expected output formats over iterative generation strategies applied to similar visual tasks (Hanu et al., 2023) leading to better accuracy (see Tab. 6.). In Section 3.3, we demonstrate both our speed gains and performance improvements.
Furthermore, Likelihood Selection is a generic method that can be easily extended to autoregressive VLMs, and in principle, there is no reason it could not also be used with extra modalities besides language. We validate this across all our experiments using the multimodal MVU framework.
Appendix G Open-Ended Video Question Answering
In this section, we explore the ability of our proposed MVU framework to operate on open-ended video question answering (QnA) tasks. For this purpose, we evaluate on the Activity-Net dataset (Yu et al., 2019) reporting the accuracy metric. We follow evaluation settings identical to Maaz et al. (2023) for these evaluations.
Given the nature of open-ended QnA tasks (i.e. no answer choices, generate free form answers), we use standard generation instead of likelihood selection. We match the generated answers against ground-truth following (Maaz et al., 2023). We present these results in Table A.5 where our MVU achieves strong results and clear improvements over the similar LLM based approach from Zhang et al. (2023a). We compare against multiple recent approaches that use similar capacity LLMs  VLMs for open-ended video QnA. We take these results as another indication to the generality of our MVU framework on video QnA tasks beyond MCQ style.
Table A.5: Open-Ended Video QnA Evaluation: We present results on the ActivityNet dataset Yu et al. (2019) that demonstrate strong performance of our proposed MVU framework. Accuracy (%) is reported. VT stands for video level training. We highlight how our MVU framework utilizes no video level training for any of its components and surpassed multiple approaches that rely on video-language training.
Method Zero-Shot VT ActivityNet-QA
JustAsk (Yang et al., 2021) 38.9
FrozenBiLM (Yang et al., 2022) 43.2
VideoCoCa (Yan et al., 2022) 56.1
FrozenBiLM (Yang et al., 2022) 24.7
Video Chat (Li et al., 2023b) 26.5
LLaMA Adapter (Zhang et al., 2023c) 34.2
Video LLaMA (Zhang et al., 2023b) 12.4
Video-ChatGPT (Maaz et al., 2023) 35.2
LocVLM (Ranasinghe et al., 2024) 37.4
Video-LLaVA (Lin et al., 2023a) 37.4
VISTA-LLaMA (Ma et al., 2023) 37.4
VideoChat-2 (Li et al., 2023c) 37.4
LLaMa-VID (Li et al., 2023e) 37.4
LLoVi (Zhang et al., 2023a) 41.8
MVU (ours) 42.2
Appendix H Additional Ablations
In this section, we repeat part of our ablation from Table 5 focused on the object motion trajectory modality inputs. We note that common video QnA benchmarks require minimal understanding of object motion to answer most questions. Our goal is to explore the value of motion information in a more relevant tasks.
Therein we investigate a new motion focused dataset, Something-Something-v2 Goyal et al. (2017) (SSv2), only for this single ablation. The SSv2 dataset focuses on motion-based category discrimination, providing an ideal evaluation to measure the usefulness of our object motion trajectory modality. We benchmark on a subset of this dataset following Lin et al. (2023b) and report these results in Table A.6. Our results while exceeding their performance also indicate the clear performance gains obtained when injecting the object motion trajectory modality into our MVU framework.
Table A.6: Ablation on Object Motion Trajectory (OMT) modality: We perform this ablation on a different dataset given the motion focused aspect we explore. Accuracy (%) reported on the motion-based SSv2 dataset clearly indicate the usefulness of the OMT modality in our MVU framework.
Method OMT Accuracy
Random - 0.6
CLIP (Radford et al., 2021) - 4.0
MAXI (Lin et al., 2023b) - 6.4
MVU (ours) 3.6
MVU (ours) 7.2
Appendix I Tokenization in LLMs
Most modern LLMs utilize Byte-Pair Encoding (BPE) tokenization Sennrich et al. (2015) to convert natural language into discrete tokens that are mapped to vocabulary embeddings. This process is learned from language itself and the resulting tokenization may sometimes break complete words into pieces (e.g. example $\rightarrow$ ex-am-ple). Given our utilization of logits extracted from an LLM forward pass, we note that each logit corresponds to a single token, which may at times be the embedding of some meaningless word piece. However, our calculation of a joint likelihood across a sequence of tokens ensures a meaningful process, which is validated by our strong results.